# Converging to Stability in Two-Sided Bandits: The Case of Unknown Preferences on Both Sides of a Matching Market

### v2

@@ A.2. Preliminaries for the Players
A.2.1. Conflict Feedback to the Players
@@
A.2.2. Arm Choice Mechanism of Players
@@
(1) $f(x)=\begin{cases}\displaystyle\frac{1-e^{-\kappa x}}{1-e^{-\kappa/2}},&0\leq x\leq 0.5,\\[4.30554pt] 1,&x>0.5,\end{cases}$
@@
A.2.3. Plausible Sets and Inconsistent Triplets
@@
That is, $S^{(i)}(t)$ is the set of all arms $a_{j}$ for which, at time-step $t-1$, some other player $p_{k}\neq p_{i}$ successfully pulled the arm, yet $p_{i}$ knows that the arm’s inherent preference favors $p_{i}$ over $p_{k}$. In CA-UCB, players had direct access to arms’ preferences, whereas in OCA-UCB, the arms’ preferences were learned deterministically. In the context of the APKP model, one may view an arm $a_{k}\in S^{(i)}(t)$ as having a win probability of 1 (with all other arms having a win probability of 0), thereby considering arms with a win probability of 1 as “plausible”. In the APU model, however, plausibility must be interpreted more generally as any arm with a non-zero probability of being won.
@@
[$\Bigl{[}\Bigl{\{}\Bigl{(}P_{ij}^{k}=1\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\leq 1-\epsilon\Bigr{)}\Bigr{\}}\vee\Bigl{\{}\Bigl{(}P_{ij}^{k}=0\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\geq\epsilon\Bigr{)}\Bigr{\}}\Bigr{]},$
@@ A.3. Main Results
A.4. Proof of Theorem 6.1
We will begin the proof of Theorem 6.1 by first establishing some important lemmas that relate to bounding the time that is taken for the arms to learn their preferences (i.e. the event $\alpha(t)$ happening). We begin by looking at the tracked reward estimates by an arm $a_{k}$ for a single pair of players $p_{i}$ and $p_{j}$.
Lemma A.0 (Single Pair of Players Overlap).
@@
Lemma A.0.1 establishes $Z$ as a sub-Gaussian random variable. Then, using the Hoeffding Inequality for Sub-Gaussian Random Variables yields:
@@
Lemma A.0 (All Pair of Players Overlap).
[…]
@@
In Lemma A.4, we have established that for an arm $a_{k}$ and any two players $p_{i}$ and $p_{j}$ s.t. $\mu_{i}^{k}>\mu_{j}^{k}$:
@@
[…]
@@
Lemma A.0 (All Arms for All Pair of Players Overlap).
@@
The proof of this lemma is a straightforward application of Lemma A.5 using the union bound over all $K$ arms and recognizing that replacing $n(t),\Delta_{\min},\text{ and }c(t)$ in the expression still maintains the upper-bound on the probability. ∎
A.4.1. Putting it together
Now, we will use Lemma A.6 to prove Theorem 6.1. Notice that the probability that we have upper-bounded in Lemma A.6 is the complement of the event $\alpha(t)$ (i.e. the confidence intervals for all the players tracked by all the arms are disjoint). Hence:
@@ A.4.1. Putting it together
A.5. Proof of Theorem 6.1
Assume that the event $\alpha$ occurs at time $T$, so that the conditions of Theorem 6.1 are satisfied. Our goal is to analyze the evolution of the players’ empirical winning probabilities, $\hat{P}_{ij}^{k}$, both before and after $\alpha(T)$.
@@
Lemma A.7 quantifies how unlikely it is for a player to have an empirical win estimate significantly below 0.5, which is crucial to prevent a player from discarding an arm due to an erroneously low conflict win estimate.
@@
Lemma A.0 (Optimistic Picking of Arm by Player).
@@
This inequality can be solved for $\kappa$ and any desired value of $\epsilon$ and $q$. Suppose then we pick a suitable $\kappa$, the player only fails to try the arm if $\hat{P}_{ij}^{k}<0.5-\epsilon$. By Lemma A.7 we know that $P(\hat{P}_{ij}^{k}<0.5-\epsilon)\leq\exp\Bigl{(}-2\,N_{c_{ij}}^{k}(T)\,\epsilon^{2}\Bigr{)}$. Therefore, with probability at least $1-\exp\Bigl{(}-2\,N_{c_{ij}}^{k}(T)\,\epsilon^{2}\Bigr{)}$, the player’s win estimate will be at least $0.5-\epsilon$, ensuring that the player continues to attempt the arm. ∎
@@
By Lemma A.7, for any players $p_{i}$ and $p_{j}$ and any arm $a_{k}$ (with $p_{i}\succ_{a_{k}}p_{j}$), we have
@@
Finally, by Lemma A.8, as long as every empirical win estimate exceeds $0.5-\epsilon$, the optimism function guarantees that players will continue selecting arms. Thus, the probability that all players continue attempting arms is at least
@@
But realize that not all conflict win estimates will converge since some triplets eventually stop receiving feedback. In the context of Theorem 6.1, before $\alpha(T)$, players are getting sample rewards as they are getting picked by the arms uniformly at random, leading to the players having some optimistic reward estimate. Consequently, after $\alpha(T)$, players have some reward estimates for the arms and a belief about conflict wins (with the worst-case estimate being around $0.5-\epsilon$). Right after $\alpha(T)$ occurs, the combined effect of the optimistic reward estimates and optimism function ensures that conflicts persist i.e. the players continue trying the arms. As more conflicts occur, the empirical win estimates gradually converge toward their true values, causing the impact of the optimism function to either be magnified (if the win probability exceeds 0.5, effectively guaranteeing a win) or diminished. This process enables players to gather more accurate information for selecting arms. In cases where a player already receives a better reward from one arm, they may cease exploring some inferior arm. This would cause the conflict win probability of the associated inferior triplet to never reach the true value. But as we will see in the following lemma, this is the case only for inconsequential triplets.
Lemma A.0.
@@
By Lemma A.10, only a bounded number of outcomes is required before $\hat{P}_{ij}^{k}$ becomes accurate enough that $(p_{i},p_{j},a_{k})$ is no longer inconsistent.
@@
Now, we are ready to complete the proof of Theorem 6.1. By design, once $\alpha$ occurs at time $T$, the feedback provided to the players is deterministic. Moreover, by Lemma A.9, the probability that the optimism function lets the players continue attempting the arms despite bad conflict win estimates is lower bounded by:
@@
where $\epsilon=\left(\Delta_{\text{min}}-2c(T)\right)\sqrt{\frac{n(T)}{8N_{c}(T)}}$ is chosen to match the tail probability in Theorem 6.1. Finally, by applying Lemma A.10, we conclude that for all consequential conflicts $(p_{i},p_{j},p_{k})$ and any $t\geq 0$, the updated empirical probability satisfies
@@
And by Lemma A.11, conflict estimates that fail to satisfy this inequality are inconsequential. This completes the proof of Theorem 6.1. ∎
@@
A.6. Proof of Theorem 6.1
The proof follows by combining Theorem 6.1, Theorem 6.1, and the supporting lemmas to show that, as $t\to\infty$, the execution of PCA-SCA converges to that of CA-UCB.
Theorem 6.1 provides a probabilistic bound on the event $\alpha(T)^{c}$ (i.e. when the arms’ confidence intervals overlap and the win feedback remains non-deterministic). Theorem 6.1 further bounds the event that the optimism function fails to motivate continued sampling, ensuring that every consequential triplet eventually receives sufficient feedback. Formally, define the failure event as things “going wrong” as:
@@
$\leq P(\text{ arms do not learn preferences })+$
$P(\text{ some triplet has a win estimate less than }0.5-\epsilon)$
[…]
where both $\delta(t)$ and $\tau(t)$ tend to zero as $t\to\infty$. Consequently, $P(\text{failure})\to 0$. Moreover, by Lemma A.11, for every consequential triplet the empirical conflict win probability converges arbitrarily close to its true value as $t\to\infty$. This implies that, in the limit, arms’ have knowledge of their own preferences, and players have knowledge of arms’ preferences (in terms of the win conflict probabilities). If we consider all the arms where the probability of winning a conflict is $\geq 1-\epsilon$ as the arms that go in the player’s plausible set, players always choose the arm with the highest UCB estimate. It then follows that, with probability tending to one, the execution of PCA-SCA becomes identical to that of CA-UCB. ∎
Appendix B Alternative Methods for Controlling Player Preference Heterogeneity
There are alternative methods for controlling heterogeneity in player preferences, for example notions like uncoordinated/coordinated markets (Ackermann et al., 2008). First, note that uncoordinated markets as defined in (Ackermann et al., 2008) essentially have uncorrelated preferences (equivalent to our model with low $\beta$), and coordinated markets introduce correlation, albeit using a different model than ours (with correlation defined on the edges of the matching graph rather than the vertices).
We ran experiments (with results averaged over 100 different runs for N=K=10) with edge-correlated preferences as defined in (Ackermann et al., 2008), the results of which are shown in Figure 5.
Figure 5. Convergence proxy for players with correlated preferences running OCA-UCB in the APKP model with $\mathcal{X}=50,\theta=90$ The line in purple is the the edge-correlation as defined in (Ackermann et al., 2008) and the rest correspond to our paper. We can see that the edge-correlated cases are comparable to when preferences are uniformly random (compare with node-correlation, where convergence time increases)
@@ Appendix B Alternative Methods for Controlling Player Preference Heterogeneity
Appendix C Visualization of the Optimism Function
Figure 6. Visualization of different levels of “optimism” for different values of $\kappa$. Note that as $\kappa$ increases, the players are more optimistic about their chances of winning a conflict if if their empirical estimate is below 0.5

### v4

@@ D.2 Preliminaries for the Players
D.2.1 Conflict Feedback to the Players
@@
D.2.2 Arm Choice Mechanism of Players
@@
$f(x)=\begin{cases}\dfrac{1-e^{-\kappa x}}{1-e^{-\kappa/2}},&0\leq x\leq 0.5,\\[4.30554pt] 1,&x>0.5\end{cases}$ (1)
@@
D.2.3 Plausible Sets and Inconsistent Triplets
@@
That is, $S^{(i)}(t)$ is the set of all arms $a_{j}$ for which, at time-step t1t-1, some other player $p_{k}\neq p_{i}$ successfully pulled the arm, yet $p_{i}$ knows that the arm’s inherent preference favors $p_{i}$ over $p_{k}$. In CA-UCB, players had direct access to arms’ preferences, whereas in OCA-UCB, the arms’ preferences were learned deterministically. In the context of the APKP model, one may view an arm $a_{k}\in S^{(i)}(t)$ as having a win probability of 1 (with all other arms having a win probability of 0), thereby considering arms with a win probability of 1 as “plausible”. In the APU model, however, plausibility must be interpreted more generally as any arm with a non-zero probability of being won.
@@
$\Bigl{[}\Bigl{\{}\Bigl{(}P_{ij}^{k}=1\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\leq 1-\epsilon\Bigr{)}\Bigr{\}}\vee\Bigl{\{}\Bigl{(}P_{ij}^{k}=0\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\geq\epsilon\Bigr{)}\Bigr{\}}\Bigr{]},$
@@ D.3 Main Results
D.4 Proof of Theorem 3
We will begin the proof of Theorem 3 by first establishing some important lemmas that relate to bounding the time that is taken for the arms to learn their preferences (i.e. the event $\alpha(t)$ happening). We begin by looking at the tracked reward estimates by an arm $a_{k}$ for a single pair of players $p_{i}$ and $p_{j}$.
Lemma D.1 (Single Pair of Players Overlap).
@@ D.4 Proof of Theorem 3
Lemma D.1.1 establishes $Z$ as a sub-Gaussian random variable. Then, using the Hoeffding Inequality for Sub-Gaussian Random Variables yields:
@@
Lemma D.2 (All Pair of Players Overlap).
Consider $a_{k}$ be a fixed arm and $\{p_{i}\}_{i=1}^{N}$ be the set of players each with true means $\{\mu_{i}^{k}\}_{i=1}^{N}$. Let $\Delta_{\min}^{k}=\min\{\mu_{i}^{k}-\mu_{j}^{k}:\mu_{i}^{k}>\mu_{j}^{k}\}$, then using $n^{k}(t)=\underset{i}{\min}\left[n_{i}^{k}(t)\right]$ and the confidence interval $c^{k}(t)=\sqrt{\frac{3\ln(t)}{2n^{k}(t)}}$, the probability that any pair of confidence intervals overlaps for arm $a_{k}$ satisfies:
@@
In Lemma D.1, we have established that for an arm $a_{k}$ and any two players $p_{i}$ and $p_{j}$ s.t. $\mu_{i}^{k}>\mu_{j}^{k}$:
@@
Notice that if we let $\Delta_{\min}^{k}=\min\{\mu_{i}^{k}-\mu_{j}^{k}:\mu_{i}^{k}>\mu_{j}^{k}\}$ and set $n_{k}(t)=\underset{i}{\min}\left[n_{i}^{k}(t)\right]$, we have that:
@@
Lemma D.3 (All Arms for All Pair of Players Overlap).
@@
The proof of this lemma is a straightforward application of Lemma D.2 using the union bound over all $K$ arms and recognizing that replacing $n(t),\Delta_{\min},\text{ and }c(t)$ in the expression still maintains the upper-bound on the probability. ∎
D.4.1 Putting it together
Now, we will use Lemma D.3 to prove Theorem 3. Notice that the probability that we have upper-bounded in Lemma D.3 is the complement of the event $\alpha(t)$ (i.e. the confidence intervals for all the players tracked by all the arms are disjoint). Hence:
@@ D.4.1 Putting it together
D.5 Proof of Theorem 3
Assume that the event $\alpha$ occurs at time $T$, so that the conditions of Theorem 3 are satisfied. Our goal is to analyze the evolution of the players’ empirical winning probabilities, $\hat{P}_{ij}^{k}$, both before and after $\alpha(T)$.
@@ D.5 Proof of Theorem 3
Lemma D.4 quantifies how unlikely it is for a player to have an empirical win estimate significantly below 0.5, which is crucial to prevent a player from discarding an arm due to an erroneously low conflict win estimate.
@@
Lemma D.5 (Optimistic Picking of Arm by Player).
@@
This inequality can be solved for $\kappa$ and any desired value of $\epsilon$ and $q$. Suppose then we pick a suitable $\kappa$, the player only fails to try the arm if $\hat{P}_{ij}^{k}<0.5-\epsilon$. By Lemma D.4 we know that $P(\hat{P}_{ij}^{k}<0.5-\epsilon)\leq\exp\Bigl{(}-2\,N_{c_{ij}}^{k}(T)\,\epsilon^{2}\Bigr{)}$. Therefore, with probability at least $1-\exp\Bigl{(}-2\,N_{c_{ij}}^{k}(T)\,\epsilon^{2}\Bigr{)}$, the player’s win estimate will be at least $0.5-\epsilon$, ensuring that the player continues to attempt the arm. ∎
@@
By Lemma D.4, for any players $p_{i}$ and $p_{j}$ and any arm $a_{k}$ (with $p_{i}\succ_{a_{k}}p_{j}$), we have
@@
Finally, by Lemma D.5, as long as every empirical win estimate exceeds $0.5-\epsilon$, the optimism function guarantees that players will continue selecting arms. Thus, the probability that all players continue attempting arms is at least
@@
But realize that not all conflict win estimates will converge since some triplets eventually stop receiving feedback. In the context of Theorem 3, before $\alpha(T)$, players are getting sample rewards as they are getting picked by the arms uniformly at random, leading to the players having some optimistic reward estimate. Consequently, after $\alpha(T)$, players have some reward estimates for the arms and a belief about conflict wins (with the worst-case estimate being around $0.5-\epsilon$). Right after $\alpha(T)$ occurs, the combined effect of the optimistic reward estimates and optimism function ensures that conflicts persist i.e. the players continue trying the arms. As more conflicts occur, the empirical win estimates gradually converge toward their true values, causing the impact of the optimism function to either be magnified (if the win probability exceeds 0.5, effectively guaranteeing a win) or diminished. This process enables players to gather more accurate information for selecting arms. In cases where a player already receives a better reward from one arm, they may cease exploring some inferior arm. This would cause the conflict win probability of the associated inferior triplet to never reach the true value. But as we will see in the following lemma, this is the case only for inconsequential triplets.
Lemma D.8.
@@
By Lemma D.7, only a bounded number of outcomes is required before $\hat{P}_{ij}^{k}$ becomes accurate enough that $(p_{i},p_{j},a_{k})$ is no longer inconsistent.
@@
Now, we are ready to complete the proof of Theorem 3. By design, once $\alpha$ occurs at time $T$, the feedback provided to the players is deterministic. Moreover, by Lemma D.6, the probability that the optimism function lets the players continue attempting the arms despite bad conflict win estimates is lower bounded by:
@@
where $\epsilon=\left(\Delta_{\text{min}}-2c(T)\right)\sqrt{\frac{n(T)}{8N_{c}(T)}}$ is chosen to match the tail probability in Theorem 3. Finally, by applying Lemma D.7, we conclude that for all consequential conflicts $(p_{i},p_{j},p_{k})$ and any $t\geq 0$, the updated empirical probability satisfies
@@
And by Lemma D.8, conflict estimates that fail to satisfy this inequality are inconsequential. This completes the proof of Theorem 3. ∎
@@
D.6 Proof of Theorem 3
The proof follows by combining Theorem 3, Theorem 3, and the supporting lemmas to show that, as $t\to\infty$, the execution of PCA-SCA converges to that of CA-UCB.
Theorem 3 provides a probabilistic bound on the event $\alpha(T)^{c}$ (i.e. when the arms’ confidence intervals overlap and the win feedback remains non-deterministic). Theorem 3 further bounds the event that the optimism function fails to motivate continued sampling, ensuring that every consequential triplet eventually receives sufficient feedback. Formally, define the failure event as things “going wrong" as:
$P(\text{ things go wrong })$ =P( arms’ confidence intervals ovelap =P($\text${ arms' confidence intervals ovelap $\cup$
[…]
$\leq P(\text{ arms do not learn preferences })+$
@@ D.6 Proof of Theorem 3
where both $\delta(t)$ and $\tau(t)$ tend to zero as $t\to\infty$. Consequently, $P(\text{failure})\to 0$. Moreover, by Lemma D.8, for every consequential triplet the empirical conflict win probability converges arbitrarily close to its true value as $t\to\infty$. This implies that, in the limit, arms’ have knowledge of their own preferences, and players have knowledge of arms’ preferences (in terms of the win conflict probabilities). If we consider all the arms where the probability of winning a conflict is $\geq 1-\epsilon$ as the arms that go in the player’s plausible set, players always choose the arm with the highest UCB estimate. It then follows that, with probability tending to one, the execution of PCA-SCA becomes identical to that of CA-UCB. ∎
Appendix E Alternative Methods for Controlling Player Preference Heterogeneity
There are alternative methods for controlling heterogeneity in player preferences, for example notions like uncoordinated/coordinated markets Ackermann et al. (2008). First, note that uncoordinated markets as defined in Ackermann et al. (2008) essentially have uncorrelated preferences (equivalent to our model with low $\beta$), and coordinated markets introduce correlation, albeit using a different model than ours (with correlation defined on the edges of the matching graph rather than the vertices).
We ran experiments (with results averaged over 100 different runs for N=K=10) with edge-correlated preferences as defined in Ackermann et al. (2008), the results of which are shown in Figure 4.
Figure 4: Convergence proxy for players with correlated preferences running OCA-UCB in the APKP model with $\mathcal{X}=50,\theta=90$ . The line in purple is the the edge-correlation as defined in Ackermann et al. (2008) and the rest correspond to our paper. We can see that the edge-correlated cases are comparable to when preferences are uniformly random (compare with node-correlation, where convergence time increases)
@@ Appendix E Alternative Methods for Controlling Player Preference Heterogeneity
Appendix F Visualization of the Optimism Function
Figure 5: Visualization of different levels of “optimism” for different values of $\kappa$. Note that as $\kappa$ increases, the players are more optimistic about their chances of winning a conflict if if their empirical estimate is below 0.5
Appendix G Machine Specifications
All experiments were run on a Apple MacBook Pro (16-inch, 2021) with the M1 Pro chip. Technical specifications:
  • Year introduced: 2021
  • Chip:
    • Apple M1 Pro: 10-core CPU (8 performance + 2 efficiency cores)
    • 16-core GPU; 16-core Neural Engine
    • 200 GB/s unified memory bandwidth
    • Media engine: HW-accelerated H.264, HEVC, ProRes & ProRes RAW; video decode/encode; ProRes encode/decode
  • Memory: 16 GB unified (on-chip) RAM
  • Storage: 512 GB SSD
  • Battery & Power:
    • 100 Wh lithium-polymer battery
Appendix H Statistical Analysis of Time to Convergence
To generate the Graphs in Figure 3, setting $\mathcal{X}=1000$ and $\theta=90$, and calculate the convergence proxy. Here, we aim to test statistically if PCA-TS converges faster than PCA-TS.We proceed as follows. For each run $j$ and market size $N$, define
$T^{(j)}=\min\bigl{\{}\,t:\text{convergence proxy}_{j}(t)=1\bigr{\}},$
and let the paired difference be:
$\Delta^{(j)}\;=\;T_{\mathrm{UCB}}^{(j)}\;-\;T_{\mathrm{TS}}^{(j)}.$
Because the distribution of $\Delta^{(j)}$ is right–skewed and contains imputed maximum values for runs that never converged within the horizon, we use two nonparametric tests:
  1. 1.
    Wilcoxon signed-rank test for $H_{0}:\ \mathrm{median}(\Delta)=0\quad\text{vs.}\quad H_{1}:\ \mathrm{median}(\Delta)>0,$ which assesses whether UCB takes longer than TS in median.
  2. 2.
    Binomial sign test for $H_{0}:\ \Pr(T_{\mathrm{TS}}<T_{\mathrm{UCB}})=0.5\quad\text{vs.}\quad H_{1}:\ \Pr(T_{\mathrm{TS}}<T_{\mathrm{UCB}})>0.5,$ which checks if TS is faster in a majority of runs.
Table 1 reports, for each market size $N$, the median difference $\mathrm{median}(\Delta)$, the Hodges–Lehmann estimator of the shift, and the $p$-values from both tests.
Table 1: Time-to-convergence summary by market size NN.
NN $\mathrm{median}(\Delta)$ Wilcoxon $p$ Binomial $p$
5 577 $5.7\times 10^{-10}$ $3.1\times 10^{-11}$
10 890 $4.6\times 10^{-9}$ $2.8\times 10^{-8}$
15 561 $9.2\times 10^{-6}$ $1.6\times 10^{-5}$
20 816 $7.7\times 10^{-4}$ $2.3\times 10^{-3}$
Results. For all tested market sizes $N$, Thompson Sampling reaches the convergence proxy of 1 significantly faster than PCA-UCB under both the Wilcoxon signed-rank test and the binomial sign test (all $p<0.01$ ). Median speed-ups range from 561 to 890 time steps.